Two waves are traveling at the same speed in a medium. One of the waves has a frequency of 256 Hz and an amplitude of 5.0 m. The other wave has a frequency of 512 Hz and an amplitude of 2.5 m. Find the ratio of their intensities.


1
Expert's answer
2021-12-20T10:30:29-0500

The ratio of intensities is


"\frac{I_1}{I_2}=\frac{A_1^2}{A_2^2}=4."